# GoMatching: A Simple Baseline for Video Text Spotting via Long and Short Term Matching

### v1

@@ Abstract
Beyond the text detection and recognition tasks in image text spotting, video text spotting presents an augmented challenge with the inclusion of tracking. While advanced end-to-end trainable methods have shown commendable performance, the pursuit of multi-task optimization may pose the risk of producing sub-optimal outcomes for individual tasks. In this paper, we highlight a main bottleneck in the state-of-the-art video text spotter: the limited recognition capability. In response to this issue, we propose to efficiently turn an off-the-shelf query-based image text spotter into a specialist on video and present a simple baseline termed GoMatching, which focuses the training efforts on tracking while maintaining strong recognition performance. To adapt the image text spotter to video datasets, we add a rescoring head to rescore each detected instance’s confidence via efficient tuning, leading to a better tracking candidate pool. Additionally, we design a long-short term matching module, termed LST-Matcher, to enhance the spotter’s tracking capability by integrating both long- and short-term matching results via Transformer. Based on the above simple designs, GoMatching achieves impressive performance on two public benchmarks, e.g., setting a new record on the ICDAR15-video dataset, and one novel test set with arbitrary-shaped text, while saving considerable training budgets. The code will be released at GoMatching.
@@ 1 Introduction
Text spotting has received increasing attention due to its various applications, such as video retrieval Dong et al. (2021) and autonomous driving Zhang et al. (2021). Recently, numerous image text spotting (ITS) methods Liu et al. (2023); Zhang et al. (2022a); Ye et al. (2023a); Huang et al. (2023) that simultaneously tackle text detection and recognition, have attained extraordinary accomplishment.
@@
Figure 1: (a) ‘Gap between Spot. & Det.’: the gap between spotting and detection F1-score. The detection score is the upper bound for spotting. The larger the gap, the poorer the recognition ability. Compared to the ITS model, the VTS model presents unsatisfactory text spotting F1-scores, which lag far behind its detection performance, especially on ArTVideo with curved text. It indicates recognition capability is a main bottleneck in the VTS model. (b) GoMatching outperforms TransDETR by over 10 MOTA on ICDAR15-video while saving 197 training GPU hours and 10.8GB memory footprint.
@@
In the video realm, video text spotting (VTS) involves a tracking task additionally. Although VTS methods Wang et al. (2017); Cheng et al. (2019, 2020); Wu et al. (2021, 2022b, 2022a) also make significant progress, a substantial discrepancy persists when compared to ITS. Specifically, excluding tracking, text spotting ability is most crucial for VTS. However, we observe that the spotting proficiency of VTS models is inferior to ITS models, especially the recognition part. To investigate this issue, we select the state-of-the-art (SOTA) VTS model Wu et al. (2022a) and ITS model Ye et al. (2023a), then compare their image-level text spotting performance on ICDAR15-video Karatzas et al. (2015) and our established ArTVideo (i.e., Arbitrary-shaped Text in Video) test set. ArTVideo is a novel test dataset established to address the scarcity of curved text in current VTS datasets. It comprises 20 high-quality annotated video clips, featuring 30% curved text approximately. As illustrated in Fig. 1(a), even when evaluated on the VTS model’s training set, the end-to-end spotting F1-score is only comparable to the zero-shot performance of the ITS model. The performance of the VTS model on ArTVideo is much worse. Moreover, there is a huge gap between the spotting and detection-only performance of the VTS model, which indicates that the recognition capability is the main bottleneck. We attribute this discrepancy to two key aspects: 1) the model architecture and 2) the training data. First, in terms of model architecture, ITS studies Ye et al. (2023a); Huang et al. (2023) have presented the advantages of employing advanced query formulation for text spotting in DETR frameworks Carion et al. (2020); Zhu et al. (2021). In contrast, Transformer-based VTS models still rely on Region of Interest (RoI) components or simply cropping detected text regions for recognition. On the other hand, Zhang et al. (2023); Yu et al. (2023) indicates that there exists optimization conflict in detection and association during end-to-end training of the MOTR Zeng et al. (2022). We hold that TransDETR Wu et al. (2022a), which further incorporates text recognition into MOTR-based architecture, may also suffer from optimization conflict. Second, regarding the training data, most text instances in current video datasets Karatzas et al. (2015); Wu et al. (2021, 2023) are straight or oriented, and the bounding box labels are only quadrilateral, which constrains the data diversity and recognition performance as well. Overall, the limitations in model architecture and data probably lead to the unsatisfactory text spotting performance of the SOTA VTS model. Hence, leveraging model and data knowledge from ITS presents considerable value for VTS.
@@
To address the aforementioned issues, a straightforward solution is to utilize an off-the-shelf SOTA image text spotter and focus the training efforts on tracking across frames, akin to tracking-by-detection methods. An important question is how to efficiently incorporate a RoI-free image text spotter for VTS. In this paper, we propose a simple baseline via lonG and short term Matching, termed GoMatching, which utilizes an off-the-shelf RoI-free image text spotter to identify text from each single frame and associates text instances across frames with a strong tracker. Specifically, we select the state-of-the-art DeepSolo Ye et al. (2023a) as the image text spotter and design a Long-Short Temporal Matching-based tracker termed LST-Matcher. Initially, to adapt the ITS model to video datasets while preserving its inherent knowledge, we introduce a rescoring mechanism. This mechanism entails training an additional lightweight text classifier called rescoring head via efficient tuning, and recalibrating confidence scores for detected instances to mitigate performance degradation caused by the image-video domain gap. The final score for each instance is determined by a fusion operation between the original score provided by the image text spotter and the calibrated score acquired from the rescoring head. The identified text instances are then sent to LST-Matcher for association. LST-Matcher first associates simple text instances between adjacent frames by a short term matching module (ST-Matcher) and then utilizes multi-frame information to assign trajectories to those difficult text instances with heavy occlusion or strong appearance changes by a long term matching module (LT-Matcher). In this manner, LST-Matcher can simultaneously encompass the advantages of both ST-Matcher and LT-Matcher. As a result, our baseline significantly surpasses existing SOTA methods by a large margin with much lower training costs, as shown in Fig. 1(b).
In summary, the contribution of this paper is four-fold:
  • We identify the limitations in current VTS methods and propose a novel and simple baseline GoMatching, which leverages an off-the-shelf SOTA image text spotter with a strong customized tracker.
  • We introduce the rescoring mechanism and long-short term matching module to adapt image text spotter to video datasets and enhance the tracker’s capabilities.
  • We establish the ArTVideo test set for addressing the absence of curved texts in current video text spotting datasets and evaluating the performance of video text spotters on videos with arbitrary-shape text.
  • Extensive experiments on public challenging datasets demonstrate the effectiveness of our baseline and its outstanding performance with less training budgets.
@@ 2.1 Multi-Object Tracking
Multi-object tracking methods follow the tracking-by-detection (TBD) or tracking-by-query-propagation (TBQP) pipeline. TBD methods Wang et al. (2020); Aharon et al. (2022); Zhang et al. (2022b) employ detectors for localization and then use association algorithms to get object trajectories. Different from extending tracks frame-by-frame, GTR Zhou et al. (2022) proposes to generate entire trajectories at once in Transformer. TBQP paradigm extends query-based object detectorsCarion et al. (2020); Zhu et al. (2021) to tracking. MOTR Zeng et al. (2022) detects object locations and serially updates its tracking queries for detecting the same items in the following frames, achieving an end-to-end solution. However, MOTR suffers from optimization conflict between detection and association Yu et al. (2023), resulting in inferior detection performance. For the VTS task which additionally involves text recognition, a naive way of training all modules end-to-end may also lead to optimization conflict. In contrast, we explore inheriting prior knowledge of text spotting from ITS models while focusing on the tracking task.
Figure 2: The overall architecture of GoMatching. Image text spotter provides text spotting results for frames. The rescoring mechanism considers both instance scores from the image text spotter and a trainable rescoring head to reduce performance degradation due to the domain gap. Long-short term matching module (LST-Matcher) assigns IDs to text instances based on the queries in long-short term frames.
@@ 2.2 Image Text Spotting
Early approaches Liao et al. (2020a); Wang et al. (2021); Liu et al. (2021) crafted RoI-based modules to bridge text detection and recognition. However, these methods ignored one vital issue, i.e., the synergy problem between the two tasks. To overcome this dilemma, recent Transformer-based methods Kittenplon et al. (2022); Liu et al. (2023); Ye et al. (2023b); Huang et al. (2023) get rid of the fetters of RoI modules, and chase a better representation for the two tasks. For example, DETR-based TESTR Zhang et al. (2022a) utilizes two decoders for each task in parallel. In contrast, DeepSolo Ye et al. (2023a) proposes a unified and explicit query form for the two tasks, without harnessing dual decoders. However, the above methods cannot perform tracking in the video.
@@ 2.3 Video Text Spotting
Compared to ITS, existing SOTA VTS methods still rely on RoI for recognition. CoText Wu et al. (2022b) adopts a lightweight text spotter with Masked-RoI, then uses several encoders to fuse features derived from the spotter, and finally feeds them to a tracking head with cosine similarity matching. TransDETR Wu et al. (2022a) performs detection and tracking under the MOTR paradigm and then uses Rotated-RoI to extract features for the subsequent recognizer. They pursue training all modules in an end-to-end manner. In comparison, we explore how to efficiently turn a RoI-free ITS model into a VTS one. We reveal the probability of freezing off-the-shelf ITS part and focusing on tracking, thereby saving training budgets while reaching SOTA performance.
@@ 3.3 Long-Short Term Matching Module
Long-short term matching module (LST-Matcher) consists of two sub-modules: the Short Term Matching module (ST-Matcher) and the Long Term Matching module (LT-Matcher), which own the same structure. ST-Matcher is steered to match simple instances between adjacent frames into trajectories, while LT-Matcher is responsible for using long term information to address the unmatched instances due to severe occlusions or strong appearance changes. Each of them contains a one-layer Transformer encoder and a one-layer Transformer decoder Zhou et al. (2022). We use a simple multi-layer perceptron (MLP) to map the filtered text instance queries into embeddings as the input, getting rid of using RoI features as in most existing MOT methods. In the encoder, historical embeddings are enhanced by self-attention. The decoder takes embeddings in the current frame as query and enhanced historical embeddings as key for cross-attention, and computes the association score matrix. The current instances are then linked to the existing trajectories composed of historical embeddings or generate new trajectories according to the association score matrix.
@@
[…]
ST-Matcher calculates a short-term trajectory-specific association score $v^{t}_{i}(e^{t}_{\hat{\alpha}^{t}_{k}},E^{S_{t}})\in\mathbb{R}^{N_{S_{t}}}$ for $i$-th instances in frame $t$, where $e^{t}_{\hat{\alpha}^{t}_{k}}\in\mathbb{R}^{D}$ is a trajectory query and $N_{S_{t}}=N_{t}+N_{t-1}$. LT-Matcher calculates a long-term trajectory-specific association score $u^{t}_{i}(e_{k},E^{L})\in\mathbb{R}^{N}$ for $i$-th instances in frame $t$, where $e_{k}\in\{e^{1}_{\hat{\alpha}^{1}_{k}},e^{2}_{\hat{\alpha}^{2}_{k}},...,e^{T}_{\hat{\alpha}^{T}_{k}}\},N=\sum^{T}_{t=1}N_{t}$. Specifically, when $v^{t}_{i}(e^{t}_{\hat{\alpha}^{t}_{k}},E^{S_{t}})=0$ and $u^{t}_{i}(e_{k},E^{L})=0$, it means no association at time $t$. Then, ST-Matcher and LT-Matcher can predict distributions of short-term and long-term associations for all instance $i$ in frame $t$ which can be written as:
@@ 3.4 Optimization
Rescoring Loss. To train the rescoring head, we following DETR Carion et al. (2020) and use Hungarian algorithm Kuhn (1955) to find a bipartite matching $\hat{\sigma}$ between the prediction set $\hat{Y}$ and the ground truth set $Y$ with minimum matching cost $\mathcal{C}$:
@@
where $\lambda_{c}$ and $\lambda_{b}$ serve as the hyper-parameters to balance different tasks. $\hat{p}_{\sigma(i)}(c_{i})$ and $\hat{b}_{\sigma(i)}$ are the probability for ground truth class $c_{i}$ and the predicition of bounding box respectively, and $b_{i}$ represents the ground truth bounding box. $\mathcal{L}_{cls}$ is the focal loss Lin et al. (2017). Specifically, the focal loss for training the rescoring head can be formulated as:
$\begin{split}\mathcal{L}_{res}=&\sum^{N}_{1}[-\mathds{1}_{\{c_{i}\neq\varnothing\}}\alpha(1-\hat{p}_{\hat{\sigma}(i)}(c_{i}))^{\gamma}\log(\hat{p}_{\hat{\sigma}(i)}(c_{i}))\\ &-\mathds{1}_{\{c_{i}=\varnothing\}}(1-\alpha)(\hat{p}_{\hat{\sigma}(i)}(c_{i}))^{\gamma}\log(1-\hat{p}_{\hat{\sigma}(i)}(c_{i}))]\end{split},$ (6)
@@
Long-Short Association Loss. In ST-Matcher, we only consider each trajectory in the universal set of adjacent frames, while in LT-Matcher we consider each trajectory in all long term frames. For each trajectory, we optimize the log-likelihood of its assignments $\hat{\alpha}_{k}$ following GTR Zhou et al. (2022):

### v2

@@ Abstract
Beyond the text detection and recognition tasks in image text spotting, video text spotting presents an augmented challenge with the inclusion of tracking. While advanced end-to-end trainable methods have shown commendable performance, the pursuit of multi-task optimization may pose the risk of producing sub-optimal outcomes for individual tasks. In this paper, we identify a main bottleneck in the state-of-the-art video text spotter: the limited recognition capability. In response to this issue, we propose to efficiently turn an off-the-shelf query-based image text spotter into a specialist on video and present a simple baseline termed GoMatching, which focuses the training efforts on tracking while maintaining strong recognition performance. To adapt the image text spotter to video datasets, we add a rescoring head to rescore each detected instance’s confidence via efficient tuning, leading to a better tracking candidate pool. Additionally, we design a long-short term matching module, termed LST-Matcher, to enhance the spotter’s tracking capability by integrating both long- and short-term matching results via Transformer. Based on the above simple designs, GoMatching delivers new records on ICDAR15-video, DSText, BOVText, and our proposed novel test with arbitrary-shaped text termed ArTVideo, which demonstrates GoMatching’s capability to accommodate general, dense, small, arbitrary-shaped, Chinese and English text scenarios while saving considerable training budgets.
@@ 1 Introduction
Text spotting has received increasing attention due to its various applications, such as video retrieval [1] and autonomous driving [2]. Recently, numerous image text spotting (ITS) methods [3, 4, 5, 6] that simultaneously tackle text detection and recognition, have attained extraordinary accomplishment. In the video realm, video text spotting (VTS) involves a tracking task additionally. Although VTS methods [7, 8, 9, 10, 11, 12] make significant progress, a substantial discrepancy persists when compared to ITS. We observe that the text recognition proficiency of VTS models is far inferior to ITS models. To investigate this issue, we compare the state-of-the-art (SOTA) VTS model TransDETR [12] and ITS model Deepsolo [5] for image-level text spotting performance on ICDAR15-video [13] and our established ArTVideo (i.e., Arbitrary-shaped Text in Video) test set (Sec.4.1) which comprises about 30% curved text.
@@
Figure 1: (a) ‘Gap between Spot. & Det.’: the gap between spotting and detection F1-score. As the spotting task involves recognizing the results of the detection process, the detection score is indeed the upper bound of spotting performance. The larger the gap, the poorer the recognition ability. Compared to the ITS model (Deepsolo [5]), the VTS model (TransDETR [12]) presents unsatisfactory image-level text spotting F1-scores, which lag far behind its detection performance, especially on ArTVideo with curved text. It indicates recognition capability is a main bottleneck in the VTS model. (b) GoMatching outperforms TransDETR by over 12 MOTA on ICDAR15-video while saving 197 training GPU hours and 10.8GB memory. Notice that since the pre-training strategies and settings vary between TransDETR and GoMatching, the comparison is focused on the fine-tuning stage.
@@
As illustrated in Fig. 1(a), even when evaluating the image-level spotting performance on the VTS model’s training set, the F1-score of TransDETR is only comparable to the zero-shot performance of Deepsolo. The performance of the VTS model on ArTVideo is much worse. Moreover, there is a huge gap between the spotting and detection-only performance of the VTS model, which indicates that the recognition capability is the main bottleneck. We attribute this discrepancy to two key aspects: 1) the model architecture and 2) the training data. First, in terms of model architecture, ITS studies [5, 6] have presented the advantages of employing advanced query formulation for text spotting in DETR frameworks [14, 15]. In contrast, existing Transformer-based VTS models still rely on Region of Interest (RoI) components or simply cropping detected text regions for recognition. On the other hand, some studies [16, 17] have indicated that there exists optimization conflict in detection and association during the end-to-end training of MOTR [18]. We hold that TransDETR [12], which further incorporates text recognition into MOTR-based architecture, may also suffer from optimization conflict. Second, regarding the training data, most text instances in current video datasets [13, 10, 19] are straight or oriented, and the bounding box labels are only quadrilateral, which constrains the data diversity and recognition performance as well. Overall, the limitations in model architecture and data probably lead to the unsatisfactory text spotting performance of the SOTA VTS model. Hence, leveraging model and data knowledge from ITS presents considerable value for VTS.
To achieve this, a straightforward approach is to take an off-the-shelf SOTA image text spotter and focus the training efforts on tracking across frames, akin to tracking-by-detection methods. An important question is how to efficiently incorporate a RoI-free image text spotter for VTS. In this paper, we propose a simple baseline via lonG and short term Matching, termed GoMatching, which leverages an off-the-shelf RoI-free image text spotter to identify text from each single frame and associates text instances across frames with a strong tracker.
@@
Specifically, we select the state-of-the-art DeepSolo [5] as the image text spotter and design a Long-Short Term Matching-based tracker termed LST-Matcher. Initially, to adapt the DeepSolo to video datasets while preserving its inherent knowledge, we freeze Deepsolo and introduce a rescoring mechanism. This mechanism entails training an additional lightweight text classifier called rescoring head via efficient tuning, and recalibrating confidence scores for detected instances to mitigate performance degradation caused by the image-video domain gap. The final score for each instance is determined by a fusion operation between the original score provided by the image text spotter and the calibrated score acquired from the rescoring head. The identified text instances are then sent to LST-Matcher for association. LST-Matcher can effectively harnesses both long- and short-term information, making it a highly capable tracker. As a result, our baseline significantly surpasses existing SOTA methods by a large margin with much lower training costs, as shown in Fig. 1(b).
In summary, the contribution of this paper is threefold. 1) We identify the limitations in current VTS methods and propose a novel and simple baseline, which leverages an off-the-shelf image text spotter with a strong customized tracker. 2) We introduce the rescoring mechanism and long-short term matching module to adapt image text spotter to video datasets and enhance the tracker’s capabilities. 3) We establish the ArTVideo test set for addressing the absence of curved texts in current video datasets and evaluating the text spotters on videos with arbitrary-shape text. Extensive experiments on public challenging datasets and the established ArTVideo test set demonstrate the effectiveness of our baseline and its outstanding performance with less training budgets. For example, GoMatching achieves the highest ranking on ICDAR15-video and DSText. Especially on bilingual dataset BOVText, GoMatching obtains a 45% improvement on MOTA compared to the recorded best performance [11]. On curved text dataset ArTVideo, GoMatching also surpasses previous SOTA method [12] by a substantial margin.
@@ 2.1 Multi-Object Tracking
Multi-object tracking methods follow the tracking-by-detection (TBD) or tracking-by-query-propagation (TBQP) pipeline. TBD methods [20, 21, 22] employ detectors for localization and then use association algorithms to get object trajectories. Different from extending tracks frame-by-frame, GTR [23] proposes to generate entire trajectories at once in Transformer. TBQP paradigm extends query-based object detectors[14, 15] to tracking. MOTR [18] detects object locations and serially updates its tracking queries for detecting the same items in the following frames, achieving an end-to-end solution. However, MOTR suffers from optimization conflict between detection and association [16, 17], resulting in inferior detection performance. For the VTS task which additionally involves text recognition, a naive way of training all modules end-to-end may also lead to optimization conflict. In contrast, we explore inheriting prior knowledge of text spotting from ITS models while focusing on the tracking task.
@@ 2.2 Image Text Spotting
Early approaches [24, 25, 26] crafted RoI-based modules to bridge text detection and recognition. However, these methods ignored one vital issue, i.e., the synergy problem between the two tasks. To overcome this dilemma, recent Transformer-based methods [27, 3, 28, 6] get rid of the fetters of RoI modules, and chase a better representation for the two tasks. For example, DETR-based TESTR [4] uses two decoders for each task in parallel. In contrast, DeepSolo [5] proposes a unified and explicit query form for the two tasks, without harnessing dual decoders. However, the above methods cannot perform tracking in the video.
@@ 2.3 Video Text Spotting
Compared to ITS, existing SOTA VTS methods still rely on RoI for recognition. CoText [11] adopts a lightweight text spotter with Masked-RoI, then uses several encoders to fuse features derived from the spotter, and finally feeds them to a tracking head with cosine similarity matching. TransDETR [12] performs detection and tracking under the MOTR paradigm and then uses Rotated-RoI to extract features for the subsequent recognizer. They pursue training all modules in an end-to-end manner. In comparison, we explore how to efficiently turn a RoI-free ITS model into a VTS one. We reveal the probability of freezing off-the-shelf ITS part and focusing on tracking, thereby saving training budgets while reaching SOTA performance.
@@ 3.1 Overview
Figure 2: The overall architecture of GoMatching. The frozen image text spotter provides text spotting results for frames. The rescoring mechanism considers both instance scores from the image text spotter and a trainable rescoring head to reduce performance degradation due to the domain gap. Long-short term matching module (LST-Matcher) assigns IDs to text instances based on the queries in long-short term frames. The yellow star sign $\bigstar$’ indicates the final output of GoMatching.
@@ 3.3 Long-Short Term Matching Module
Long-short term matching module (LST-Matcher) consists of two sub-modules: the Short Term Matching module (ST-Matcher) and the Long Term Matching module (LT-Matcher), which own the same structure. ST-Matcher is steered to match simple instances between adjacent frames into trajectories, while LT-Matcher is responsible for using long term information to address the unmatched instances due to severe occlusions or strong appearance changes. Each of them contains a one-layer Transformer encoder and a one-layer Transformer decoder [23]. We use a simple multi-layer perceptron (MLP) to map the filtered text instance queries into embeddings as the input, getting rid of using RoI features as in most existing MOT methods. In the encoder, historical embeddings are enhanced by self-attention. The decoder takes embeddings in the current frame as query and enhanced historical embeddings as key for cross-attention, and computes the association score matrix. The current instances are then linked to the existing trajectories composed of historical embeddings or generate new trajectories according to the association score matrix.
@@
$\hat{\alpha}^{t}_{k}=\left\{\begin{array}[]{ll}\emptyset,\quad\text{if }\tau^{t}_{k}=\emptyset\text{ or max}_{i}(\textit{IoU}(b^{t}_{i},\tau^{t}_{k}))<0.5\\ \text{argmax}_{i}(\textit{IoU}(b^{t}_{i},\tau^{t}_{k})),\quad\text{otherwise}\end{array}.\right.$ (1)
ST-Matcher calculates a short-term association score $v^{t}_{i}(e^{t}_{\hat{\alpha}^{t}_{k}},E^{S_{t}})\in\mathbb{R}^{N_{S_{t}}}$ for $i$-th instances in frame $t$, where $e^{t}_{\hat{\alpha}^{t}_{k}}\in\mathbb{R}^{D}$ is a trajectory query and $N_{S_{t}}=N_{t}+N_{t-1}$. LT-Matcher calculates a long-term trajectory-specific association score $u^{t}_{i}(e_{k},E^{L})\in\mathbb{R}^{N}$ for $i$-th instances in frame $t$, where $e_{k}\in\{e^{1}_{\hat{\alpha}^{1}_{k}},e^{2}_{\hat{\alpha}^{2}_{k}},...,e^{T}_{\hat{\alpha}^{T}_{k}}\},N=\sum^{T}_{t=1}N_{t}$. Specifically, when $v^{t}_{i}(e^{t}_{\hat{\alpha}^{t}_{k}},E^{S_{t}})=0$ and $u^{t}_{i}(e_{k},E^{L})=0$, it means no association at time $t$. Then, ST-Matcher and LT-Matcher can predict distributions of short-term and long-term associations for all instance $i$ in frame $t$ which can be written as:
@@ 3.4 Optimization
Rescoring Loss. To train the rescoring head, we following DETR [14] and use Hungarian algorithm [29] to find a bipartite matching $\hat{\sigma}$ between the prediction set $\hat{Y}$ and the ground truth set $Y$ with minimum matching cost $\mathcal{C}$:
@@
where $\lambda_{c}$ and $\lambda_{b}$ serve as the hyper-parameters to balance different tasks. $\hat{p}_{\sigma(i)}(c_{i})$ and $\hat{b}_{\sigma(i)}$ are the probability for ground truth class $c_{i}$ and the predicition of bounding box respectively, and $b_{i}$ represents the ground truth bounding box. $\mathcal{L}_{cls}$ is the focal loss [30]. Specifically, the focal loss for training the rescoring head can be formulated as:
$\mathcal{L}_{res}=\sum^{N}_{1}[-\mathds{1}_{\{c_{i}\neq\varnothing\}}\alpha(1-\hat{p}_{\hat{\sigma}(i)}(c_{i}))^{\gamma}\log(\hat{p}_{\hat{\sigma}(i)}(c_{i}))-\mathds{1}_{\{c_{i}=\varnothing\}}(1-\alpha)(\hat{p}_{\hat{\sigma}(i)}(c_{i}))^{\gamma}\log(1-\hat{p}_{\hat{\sigma}(i)}(c_{i}))],$ (6)
@@
Long-Short Association Loss. In ST-Matcher, we only consider each trajectory in the universal set of adjacent frames, while in LT-Matcher we consider each trajectory in all long term frames. For each trajectory, we optimize the log-likelihood of its assignments $\hat{\alpha}_{k}$ following GTR [23]: